# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v1

@@ B.2 Closed-Source Models
Among closed-source ALLMs, Google’s Gemini series [66] and OpenAI’s GPT-4o series [54] represent the industry’s state-of-the-art in audio understanding and interaction technologies. In our evaluation of various safety concerns, we employ both the Gemini and GPT-4o model series.
Gemini-1.5 Pro leverages a Mixture-of-Experts architecture for unified reasoning across speech, image, and text. It supports audio inputs up to 19 hours in duration and contexts up to the million-token scale, enabling seamless processing for tasks such as audio summarization, transcription, and translation.
Gemini-2.5 Flash retains the core multimodal design of the Pro version while significantly optimizing inference speed and computational efficiency. This version supports up to 8.4 hours of audio input and million-token context windows, with dramatically reduced latency and operational cost compared to the Pro variant, while still covering tasks like audio summarization, transcription, and translation.
Gemini-2.5 Pro further advances multimodal reasoning, introducing a dynamic “thinking budget” mechanism that adaptively allocates computational resources based on instruction and system constraints. Its superior performance on video understanding benchmarks extends to the audio domain, enabling streaming responses for complex tasks such as conversational QA, scenario retrieval, and reasoning through efficient temporal alignment and cross-modal integration.
GPT-4o Audio is the first developer-oriented interactive audio model that supports both understanding and generation of speech. It is capable of speech transcription, summarization, sentiment analysis, and conversational dialogue.
GPT-4o mini Audio is designed to deliver cost-effective yet robust audio understanding and generation. It supports a variety of audio input formats and can produce seamless bimodal (text and speech) output with customizable speech styles, making it applicable to edge devices and large-scale embedded deployments.
@@ Appendix C Platform Design of AudioTrust
The inference execution stage focuses on raw data processing and the collection of model outputs. First, the data engine module efficiently loads and preprocesses various standard trustworthiness benchmark datasets, including both publicly released open benchmarks and custom-built datasets, thus ensuring data consistency and traceability. Subsequently, users can flexibly specify evaluation models, datasets, evaluation targets, and runtime parameters through configuration files. This enables batch parallel scheduling and significantly optimizes computational resource usage. The core Inference Module supports mainstream Audio LLM inference tasks, allowing direct loading of open-source models from the Hugging Face Hub, and natively integrates adapters for closed-source models accessed via APIs, thereby providing comprehensive full-stack support for major ALLMs. Through the aforementioned workflow, structured raw model output files are generated for subsequent analysis, ensuring a highly reproducible evaluation process.
@@ D.1 Dataset Classification Criteria
We utilized seven sensitive attributes to ensure both group and individual fairness: gender, race, age, accent, socioeconomic status, personality traits, and speech fluency. All sensitive attributes were defined with binary values. Specifically, the sensitive attribute sets were as follows: gender , age , race , accent , linguistic characteristic , economic status , and personality traits .
@@ D.2 Dataset Construction Method
Multifaceted social biases: We consider common societal biases associated with multiple sensitive attributes such as gender, race, and age [13, 101, 70, 25]. In addition to these attributes, we include distinctive stereotypes uniquely identifiable through audio characteristics, including accent, personality traits, economic status, and linguistic characteristic.
@@
During the data construction phase, we utilized the state-of-the-art multimodal generation model GPT-4o [54] to produce textual content, which was then transformed into audio using the F5-TTS model [12]. The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to ensure natural language fluency and semantic consistency. To enable controlled experiments with single-variable variation, each data pair consists of two audio samples with identical semantic content, differing solely in their associated sensitive attributes.
For example, to simulate racial identity, we altered the input language to represent different ethnic groups. GPT-4o was employed to ensure that the translated texts conveyed semantically equivalent meanings across languages, thereby eliminating confounding effects introduced by lexical or syntactic discrepancies. For sensitive attributes that cannot be directly simulated by changing the voice source—such as linguistic style or socioeconomic status—we implemented more sophisticated data augmentation strategies.
To simulate socioeconomic status, we enriched the base audio with representative background sounds—such as upscale restaurant ambience for high-income individuals and street market noise for lower-income individuals—to enhance the perceptual cues of social identity. Regarding linguistic style, we adjusted the fluency and structure of the input text before feeding it into the TTS model. This allowed us to generate speech samples that reflect varying levels of coherence, prosody, and articulation, simulating differences in communication style associated with social or educational background.
@@ D.3 Evaluation Method and Metrics
We utilized statistical methods and developed the unfairness metrics to quantitatively assess model fairness [81, 15]. Let an audio–language model be , mapping the audio–text input space to the text output space . Denote by the set of sensitive attribute values (e.g. ). For a test collection , we introduce a discriminator detecting sensitive attributes in generations.
@@
where each probability is approximated via -sample Monte-Carlo estimates. A positive implies a fairness towards group . When aggregating across tasks, we report the absolute value to emphasise fairness magnitude only.
@@ D.4 Experimental Design and Results
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | ||
@@
| Female | -1.00 | -1.00 | 1.00 | 0.90 | -0.50 | 0.58 | 0.85 | 1.00 | 0.65 | |
| Old | -0.70 | -1.00 | 1.00 | 0.00 | 0.65 | 0.47 | 1.00 | 0.50 | 0.60 | |
| American | -0.40 | -0.70 | 1.00 | 0.00 | 0.70 | 1.00 | 1.00 | 0.90 | 0.50 | |
| Clam | -0.20 | -1.00 | 1.00 | 0.45 | 0.50 | 0.80 | 0.70 | 1.00 | 1.00 | |
| Fluent | -1.00 | -0.90 | 1.00 | 0.35 | 0.90 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Chinese | 0.30 | -0.60 | 1.00 | 1.00 | 0.50 | 0.26 | 0.30 | 0.00 | 0.00 | |
| Wealthy | 1.00 | -0.60 | 1.00 | 0.87 | 1.00 | 0.58 | -0.90 | 0.20 | 0.90 | |
@@
| Female | -0.95 | -0.10 | 1.00 | 0.90 | 0.65 | 0.80 | 1.00 | 0.80 | 0.70 | |
| Old | -1.00 | -0.10 | 1.00 | 0.50 | 0.10 | 0.50 | -0.50 | 0.70 | 0.90 | |
| American | -0.70 | 0.90 | 1.00 | 1.00 | 0.60 | 0.50 | 0.50 | 0.50 | 0.90 | |
| Clam | -0.75 | -0.10 | 1.00 | 1.00 | 0.40 | 0.70 | -0.30 | 1.00 | 1.00 | |
| Fluent | 0.40 | -1.00 | 1.00 | 1.00 | 0.80 | 0.80 | 1.00 | 0.90 | 0.80 | |
| Chinese | 0.00 | 0.90 | 1.00 | 1.00 | -0.75 | 0.89 | 1.00 | 0.75 | 0.80 | |
| Wealthy | 0.85 | 0.90 | 1.00 | 0.20 | -0.10 | 0.80 | 1.00 | 0.80 | 0.50 | |
@@
| Female | 0.21 | 0.50 | 0.05 | 0.90 | 0.00 | 1.00 | 1.00 | 0.80 | 0.60 | |
| Old | -0.15 | 0.50 | 0.90 | 0.05 | 0.10 | 0.89 | 0.90 | 0.65 | 0.90 | |
| American | 0.57 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 0.90 | |
| Clam | -0.05 | -0.20 | 1.00 | 1.00 | 0.50 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Fluent | 0.67 | 0.90 | 0.20 | 0.55 | 0.60 | 0.70 | 0.90 | 1.00 | 1.00 | |
| Chinese | 0.58 | 1.00 | 1.00 | 0.60 | 0.20 | 0.68 | 0.75 | 0.05 | 0.20 | |
| Wealthy | 0.80 | 1.00 | 1.00 | 0.00 | 0.80 | 1.00 | -0.10 | 0.70 | 1.00 | |
| Average | 0.59 | 0.71 | 0.91 | 0.63 | 0.54 | 0.76 | 0.80 | 0.73 | 0.76 | |
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | ||
@@
| Female | 1.00 | 0.58 | 1.00 | 0.50 | -1.00 | -0.20 | 0.00 | -0.20 | 0.20 | |
| Old | 0.90 | 0.35 | 1.00 | 0.00 | 1.00 | 0.89 | 0.90 | 0.60 | 0.65 | |
| American | 0.90 | 0.25 | 1.00 | 0.20 | -0.40 | 0.70 | 0.30 | 0.20 | 0.90 | |
| Clam | 1.00 | 0.70 | 1.00 | -0.30 | 0.50 | 0.68 | 0.90 | 0.00 | 0.16 | |
| Fluent | 0.10 | 0.39 | 1.00 | 0.60 | 1.00 | 0.00 | 0.60 | 0.00 | 0.35 | |
| Chinese | 0.30 | 0.00 | 1.00 | -0.60 | 1.00 | 0.20 | 0.10 | 0.15 | 0.10 | |
| Wealthy | -0.30 | 0.90 | 1.00 | -0.30 | -0.30 | -0.60 | -0.80 | 0.05 | 0.25 | |
@@
| Female | 1.00 | 1.00 | 0.60 | 1.00 | 0.30 | -0.20 | 0.45 | 0.00 | 0.00 | |
| Old | 1.00 | 1.00 | 0.38 | 1.00 | 1.00 | 1.00 | 0.80 | 0.00 | 0.00 | |
| American | 1.00 | 1.00 | 0.00 | 1.00 | 0.80 | 0.40 | 0.95 | 0.00 | 0.10 | |
| Clam | 1.00 | 1.00 | 0.90 | 1.00 | 1.00 | 1.00 | 1.00 | 0.15 | 0.05 | |
| Fluent | 1.00 | 1.00 | 0.21 | 1.00 | 0.90 | 0.75 | 0.90 | 0.05 | 0.00 | |
| Chinese | 1.00 | 0.54 | 1.00 | 1.00 | -0.33 | 0.75 | 0.75 | 0.00 | 0.00 | |
| Wealthy | 1.00 | 1.00 | 1.00 | 1.00 | 0.50 | -0.79 | -0.80 | 0.10 | 0.00 | |
@@
| Female | -0.10 | 0.60 | 1.00 | -0.90 | -0.33 | -0.89 | -0.50 | 0.00 | 0.00 | |
| Old | 1.00 | 0.20 | 1.00 | -1.00 | 0.44 | 0.78 | 1.00 | 0.00 | 0.00 | |
| American | 0.70 | 0.00 | 1.00 | -0.70 | -1.00 | 0.50 | 0.40 | 0.00 | 0.10 | |
| Clam | -0.30 | 0.90 | 1.00 | -1.00 | 0.10 | 1.00 | 0.95 | 0.05 | 0.00 | |
| Fluent | -0.30 | 0.90 | 1.00 | -0.90 | 0.86 | 0.42 | 0.50 | 0.00 | 0.00 | |
| Chinese | 1.00 | -0.70 | 1.00 | -1.00 | -1.00 | 0.70 | 0.70 | 0.00 | 0.00 | |
| Wealthy | 0.90 | -1.00 | 1.00 | -1.00 | -1.00 | -0.78 | -1.00 | 0.00 | 0.00 | |
| Average | 0.75 | 0.67 | 0.86 | 0.76 | 0.70 | 0.63 | 0.68 | 0.07 | 0.14 | |
@@ Results.
We evaluate the social decision-making fairness and stereotype fairness of nine models in Appendix D.4, Table 7, Table 8, Table D.4. The results reveal the following key findings:
@@
(1) The fairness levels vary significantly among different ALLMs. Notably, models generally considered highly capable, such as GPT-4o Audio, GPT-4o mini Audio, Gemini-2.5 Flash, and Gemini-2.5 Pro, exhibit the highest group unfairness in the decision-making experiments. In contrast, some lower-performing open-source models, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox, demonstrate relatively better fairness. However, these models still exhibit high group unfairness and are far from ideal models. (2) Overall, the model’s responses tend to favor sensitive attributes such as female, old, American accent, calm, fluent, Western, and wealthy. (3) In the stereotype experiments, GPT-4o Audio and GPT-4o mini Audio show excellent fairness, while MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox exhibit the highest unfairness. Interestingly, GPT-4o Audio and GPT-4o mini Audio perform well in stereotype experiments by almost refusing to answer all harmful questions (the proportion of responses across attributes is nearly consistent), but do not refuse in decision-making tasks. This indicates that the GPT series models face challenges in accurately determining whether a question is genuinely harmful. (4) Most models that exhibit high group (un)fairness when evaluating social stereotypes, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, Ultravox, and the Gemini series, also maintain similar levels in decision-making scenarios. (5) ALLMs exhibit basically the same degree of unfairness across different scenarios. (6) The performance of LLMs in decision-making scenarios is worse compared to ALLMs.
@@ E.2 Dataset Construction Method
To construct the datasets for physical and logical violations, we adopted a two-stage procedure. First, we utilized GPT-4o [54] and audio data from Freesound111http://freesound.org/ to generate 80 sounds that represent scenarios with physical or temporal logical inconsistencies. Subsequently, we edited the collected audio content and arranged and concatenated them according to the scenarios generated by GPT-4o. The choice of GPT-4o for scene generation is attributed to its advanced linguistic capabilities and alignment guarantees, which ensure both the diversity and reliability of the generated scenarios.
In addition, to create audio corresponding to content mismatches and label mismatches, we incorporated the emotional speech dataset [105] and obtained music classification datasets from Freesound that align with our testing objectives. To expose these vulnerabilities, we randomly associated mismatched emotion labels with the audio. To ensure controllable model outputs and the reliability of the evaluation metrics, we opted to randomly recombine audio and text classification labels without altering the classification types of the original datasets.

### v3

@@ B.2 Closed-Source Models
Among closed-source ALLMs, Google’s Gemini series (Team et al., 2023) and OpenAI’s GPT-4o series (OpenAI et al., 2024) represent the industry’s state-of-the-art in audio understanding and interaction technologies. In our evaluation of various safety concerns, we employ both the Gemini and GPT-4o model series.
10. Gemini-1.5 Pro leverages a Mixture-of-Experts architecture for unified reasoning across speech, image, and text. It supports audio inputs up to 19 hours in duration and contexts up to the million-token scale, enabling seamless processing for tasks such as audio summarization, transcription, and translation.
11. GPT-4o Audio is the first developer-oriented interactive audio model that supports both understanding and generation of speech. It is capable of speech transcription, summarization, sentiment analysis, and conversational dialogue.
12. GPT-4o mini Audio is designed to deliver cost-effective yet robust audio understanding and generation. It supports a variety of audio input formats and can produce seamless bimodal (text and speech) output with customizable speech styles, making it applicable to edge devices and large-scale embedded deployments.
13. Gemini-2.5 Flash retains the core multimodal design of the Pro version while significantly optimizing inference speed and computational efficiency. This version supports up to 8.4 hours of audio input and million-token context windows, with dramatically reduced latency and operational cost compared to the Pro variant, while still covering tasks like audio summarization, transcription, and translation.
14. Gemini-2.5 Pro further advances multimodal reasoning, introducing a dynamic “thinking budget” mechanism that adaptively allocates computational resources based on instruction and system constraints. Its superior performance on video understanding benchmarks extends to the audio domain, enabling streaming responses for complex tasks such as conversational QA, scenario retrieval, and reasoning through efficient temporal alignment and cross-modal integration.
@@ Appendix C Platform Design of AudioTrust
The inference execution stage focuses on raw data processing and the collection of model outputs. First, the data engine module efficiently loads and preprocesses various standard trustworthiness benchmark datasets, including both publicly released open benchmarks and custom-built datasets, thus ensuring data consistency and traceability. Subsequently, users can flexibly specify evaluation models, datasets, evaluation targets, and runtime parameters through configuration files. This enables batch parallel scheduling and significantly optimizes computational resource usage. The core inference module supports mainstream ALLMs inference tasks, allowing direct loading of open-source models from the Hugging Face Hub, and natively integrates adapters for closed-source models accessed via APIs, thereby providing comprehensive full-stack support for major ALLMs. Through the aforementioned workflow, structured raw model output files are generated for subsequent analysis, ensuring a highly reproducible evaluation process.
@@
Platform Validation.
Beyond automated evaluators, we include a human-aided validation protocol: 10% of the data are randomly sampled and cross-checked by 20 trained annotators, with each QA pair independently reviewed by three annotators. Final labels are decided by majority voting, yielding a 97–98% agreement with the platform’s automated assessments, thereby validating the reliability of the framework.
@@ D.1 Dataset Classification Criteria
We utilized seven sensitive attributes to ensure both group and individual fairness: gender, race, age, accent, economic status, personality traits, and speech fluency. All sensitive attributes were defined with binary values. Specifically, the sensitive attribute sets were as follows: gender , age , race , accent , linguistic characteristic , economic status , and personality traits .
@@ D.2 Dataset Construction Method
Multifaceted social biases: We consider common societal biases associated with multiple sensitive attributes such as gender, race, and age (Cho et al., 2023; Zhang et al., 2017b; Wan et al., 2024; Hirota et al., 2023). In addition to these attributes, we include distinctive stereotypes uniquely identifiable through audio characteristics, including accent, personality traits, economic status, and linguistic characteristic.
@@
During the data construction phase, we utilized the multi-modal generation model GPT-4o (OpenAI et al., 2024) to produce textual content, which was then transformed into audio using the F5-TTS model (Chen et al., 2024). The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to ensure natural language fluency and semantic consistency. To enable controlled experiments with single-variable variation, each data pair consists of two audio samples with identical semantic content, differing solely in their associated sensitive attributes.
For example, to simulate racial identity, we altered the input language to represent different ethnic groups. GPT-4o was employed to ensure that the translated texts conveyed semantically equivalent meanings across languages, thereby eliminating confounding effects introduced by lexical or syntactic discrepancies. For sensitive attributes that cannot be directly simulated by changing the voice source, such as linguistic style or economic status, we implemented more sophisticated data augmentation strategies.
To simulate economic status, we enriched the base audio with representative background sounds to enhance the perceptual cues of social identity. Examples include upscale restaurant ambience for high-income individuals and street market noise for lower-income individuals. individuals. Regarding linguistic style, we adjusted the fluency and structure of the input text before feeding it into the TTS model. This allowed us to generate speech samples that reflect varying levels of coherence, prosody, and articulation, simulating differences in communication style associated with social or educational background.
@@ D.3 Evaluation Method and Metrics
We utilized statistical methods and developed the unfairness metrics to quantitatively assess model fairness (Xu et al., 2025a; Deldjoo & Nazary, 2024). Let an audio–language model be , mapping the audio–text input space to the text output space . Denote by the set of sensitive attribute values (e.g. ). For a test collection , we introduce a discriminator detecting sensitive attributes in generations.
@@
where each probability is approximated via ‑sample Monte‑Carlo estimates. A positive implies a fairness towards group . When aggregating across tasks, we report the absolute value to emphasise fairness magnitude only.
@@ D.4 Experimental Design and Results
| SALMONN | Ultravox | Qwen2-Audio | MiniCPM-o 2.6 | Step-Fun | Qwen2.5-omni | Kimi-Audio | OpenS2S | Step-Audio2 | ||
@@
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
@@
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
@@
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Average | 0.911 | 0.631 | 0.710 | 0.585 | 0.495 | 0.798 | 0.914 | 0.843 | 0.902 | |
| Gemini-1.5 Pro | GPT-4o Audio | GPT-4o mini Audio | Gemini-2.5 Flash | Gemini-2.5 Pro | ||
| Recruitment | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Admission | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Financial Loan | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Average | 0.540 | 0.726 | 0.755 | 0.759 | 0.795 | |
| SALMONN | Ultravox | Qwen2-Audio | MiniCPM-o 2.6 | Step-Fun | Qwen2.5-omni | Kimi-Audio | OpenS2S | Step-Audio2 | ||
@@
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
@@
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
@@
| Female | ||||||||||
| Old | ||||||||||
| American | ||||||||||
| Clam | ||||||||||
| Fluent | ||||||||||
| Chinese | ||||||||||
| Wealthy | ||||||||||
| Average | 0.861 | 0.762 | 0.667 | 0.740 | 0.342 | 0.933 | 0.964 | 0.983 | 0.926 | |
| Gemini-1.5 Pro | GPT-4o Audio | GPT-4o mini Audio | Gemini-2.5 Flash | Gemini-2.5 Pro | ||
| Occupational | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Education | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Medical | ||||||
| Female | ||||||
| Old | ||||||
| American | ||||||
| Clam | ||||||
| Fluent | ||||||
| Chinese | ||||||
| Wealthy | ||||||
| Average | 0.703 | 0.070 | 0.140 | 0.630 | 0.680 | |
@@ Results.
We evaluate the social decision-making fairness and stereotype fairness of nine models in Appendix D.4, Table 7, Table 8, Table 10, Table 9, Table D.4. The results reveal the following key findings:
@@
(1) The fairness levels vary significantly among different ALLMs. Notably, models generally considered highly capable, such as GPT-4o Audio, GPT-4o mini Audio, Gemini-2.5 Flash, and Gemini-2.5 Pro, exhibit the highest group unfairness in the decision-making experiments. In contrast, some lower-performing open-source models, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox, and Step-Fun, demonstrate relatively better fairness. However, these models still exhibit high group unfairness and are far from ideal models. (2) Overall, the model’s responses tend to favor sensitive attributes such as female, old, American accent, calm, fluent, Western, and wealthy. (3) In the stereotype experiments, GPT-4o Audio and GPT-4o mini Audio show excellent fairness, while MiniCPM-o 2.6, Qwen2-Audio, SALMONN, and Ultravox exhibit the highest unfairness. Interestingly, GPT-4o Audio and GPT-4o mini Audio perform well in stereotype experiments by almost refusing to answer all harmful questions (the proportion of responses across attributes is nearly consistent), but do not refuse in decision-making tasks. This indicates that the GPT series models face challenges in accurately determining whether a question is genuinely harmful. (4) Most models that exhibit high group (un)fairness when evaluating social stereotypes, such as MiniCPM-o 2.6, Qwen2-Audio, SALMONN, Ultravox, and the Gemini series, also maintain similar levels in decision-making scenarios. (5) ALLMs exhibit basically the same degree of unfairness across different scenarios. (6) The performance of LLMs in decision-making scenarios is worse compared to ALLMs.
@@ E.2 Dataset Construction Method
To construct the datasets for physical and logical violations, we adopted a two-stage procedure. First, we utilized GPT-4o (OpenAI et al., 2024) and audio data from Freesound111http://freesound.org/ to generate 80 sounds that represent scenarios with physical or temporal logical inconsistencies. Subsequently, we edited the collected audio content and arranged and concatenated them according to the scenarios generated by GPT-4o. The choice of GPT-4o for scene generation is attributed to its advanced linguistic capabilities and alignment guarantees, which ensure both the diversity and reliability of the generated scenarios.
In addition, to create audio corresponding to content mismatches and label mismatches, we incorporated the emotional speech dataset (Zhou et al., 2022) and obtained music classification datasets from Freesound that align with our testing objectives. To expose these vulnerabilities, we randomly associated mismatched emotion labels with the audio. To ensure controllable model outputs and the reliability of the evaluation metrics, we opted to randomly recombine audio and text classification labels without altering the classification types of the original datasets.